# On the Robustness of Adversarial Training Against Uncertainty Attacks

### v1

@@ 4 Theoretical Analysis on Adversarial Uncertainty Robustness
A recent work [7] likewise raised this observation, but it is still limited to empirically analyzing a small set of models and only against specific attack instances. To this aim, instead of considering specific attack instances, we answer our research question by analyzing the Uncertainty Span of adversarial trained models, i.e., consider both the under- and over-confidence cases, in accordance with  Sect. 3, drawing general conclusions on these models’ robustness against any attack instance. Since the predictive entropy of a sample increases when approaching the decision boundary, it is reasonable to think that adversarial training may also protect against under-confidence attacks: in fact, adversarial training forces the model to be robust to label-flips (i.e., to samples crossing the boundary), hindering any potential attacker interested in increasing uncertainty. For over-confidence attacks instead, we cannot draw similar conclusions: indeed, being robust to label flips does not guarantee an entropy reduction since the regions where the entropy is minimized do not lie on the decision boundary.
@@
4.1 On the Convergence of Adversarial Trained Models
@@
Lemma 4.1 represents a well-known result of probabilistic machine learning [15], stating that the optimal classifier should be perfectly calibrated. This phenomenon should be considered when dealing with aleatoric uncertainty because it asserts that lower entropy does not always go along with more accurate classifiers. From this starting point, we will now show that such conditions slightly change in an attacker’s presence.
@@
$\mathcal{L}^{\boldsymbol{\theta}_{R}}(\mathcal{D}_{\textrm{adv}}^{\boldsymbol{x}})$ $=\underbrace{p(\textrm{A}|\boldsymbol{x}+\delta_{\textrm{A}})}_{z}\cdot\underbrace{H(\boldsymbol{t_{A}},f^{\boldsymbol{\theta}}(\boldsymbol{x}+\delta_{\textrm{A}}))}_{-\log(\alpha-\beta_{A})}+\underbrace{p(\textrm{B}|\boldsymbol{x}+\delta_{\textrm{B}})}_{1-z}\cdot\underbrace{H(\boldsymbol{t_{B}},f^{\boldsymbol{\theta}}(\boldsymbol{x}+\delta_{\textrm{B}}))}_{-\log(1-\alpha-\beta_{B})}$
@@
Contrary to Lemma 4.1, when using adversarial training, the loss does not equal a cross entropy between two distinct components. Therefore, we need to compute the values for which the partial derivative with respect to $\alpha$ equals zero:
$\implies\frac{\partial\mathcal{L}}{\partial\alpha}=0\implies\frac{1-z}{1-\alpha-\beta}-\frac{z}{\alpha-\beta}=0\implies\alpha=z-\beta\cdot(2z-1)$
4.2 Demystifying the Underconfidence of Robust Models
From Lemma 4.1 and Lemma 4.2, it is possible to derive that in the presence of an attacker, a classifier inevitably deviates the trajectory of the optimization process from the one it would follow in a standard, clean optimization process. Interestingly, the entropy of an adversarial-trained classifier is higher compared to standard training. We can easily prove this inequality starting from the two lemmas:
Theorem 4.1.
@@ 4.2 Demystifying the Underconfidence of Robust Models
From Lemma 4.1, we know $f^{\boldsymbol{\theta}_{S}}(x)=[z,1-z]^{\intercal}$, while from Lemma 4.2 we know that in the presence of an attacker, the optimum deviates from an offset amount of $o=\beta(2z-1)$: $f^{\boldsymbol{\theta}_{R}}(x)=[z-o,1-z+o]^{\intercal}$. Within this interval, the entropy $H(f^{\boldsymbol{\theta}_{R}}(x))$ reaches its minimum when the offset equals zero and, thus, when $f^{\boldsymbol{\theta}_{S}}(x)=f^{\boldsymbol{\theta}_{R}}(x)$; therefore, $H(f^{\boldsymbol{\theta}_{S}}(x))\leq H(f^{\boldsymbol{\theta}_{R}}(x))$. ∎
@@
From Theorem 4.1, we can draw some conclusions on the robustness against uncertainty attacks by studying the effect of adversarial training on the Uncertainty Span. Albeit  Theorem 4.1 proves the tendency of adversarial training to produce underconfident models, it remains factual the robustness of such models against traditional adversarial examples, i.e., crafted to induce a misclassification on a sample [11]. In order to succeed, the adversarial example requires crossing the decision boundary; in a binary classification problem, such conditions correspond precisely to reaching the maximum possible entropy. Consequently, reducing the number of samples for which a traditional adversarial example exists (i.e., increasing the robust accuracy) directly implies reducing the upper bound of the uncertainty span. Concerning the robustness against over-confidence attacks, findings from  Theorem 4.1 suggest greater robustness against such attacks as well. This theorem shows that adversarial-trained models output predictions with higher entropies than undefended models. Suppose we aim to decrease it until reaching a desired value of $v$: in that case, the initially higher entropy of the adversarial trained model makes it more challenging to lower it to $v$, as the starting values are already higher and, therefore, need a significant reduction.
@@
Discussion – In safe environments (without any attacker), a model whose purpose is to solve a classification problem takes advantage of modeling correctly the true posterior $p(y|\boldsymbol{x})$ for any given input sample. Nevertheless, in the presence of an attacker, the classifier tends to balance good generalization capabilities and security guarantees; such a balance compromises not only the model’s accuracy (as we already know from previous literature) but also comes at the price of sacrificing a certain amount of confidence in the predictions. As a result, adversarial-trained models disclose robust characteristics against uncertainty attacks, providing theoretical support for the claim that adversarial training makes models more resilient to over- and under-confidence attacks.
5 Experimental Analysis
@@
We now present the experimental setup employed to empirically validate our theoretical findings by analysing the Uncertainty Span on a selection of 23 state-of-the-art models from the well-known Robustbench repository [12]. In addition, we also consider a concrete case study where an attacker is interested in tampering with the model’s calibration, proving the robustness of adversarial trained models for a real scenario. The code for the reproducibility of the experiments is available at the following link: https://github.com/EmanueleLedda97/UncertaintyAdversarialRobustness
5.1 Experimental Setup
@@
Models – For a comprehensive investigation, we consider 15 CIFAR-10 (denoted as C1-C15) and 8 ImageNet (denoted as I1-I8) standard robust classifiers from the Robustbench repository [12], encompassing many different model architectures and robust accuracies. Many of them rely on data augmentation (C2 [18], C3, C4 [19], C5,C6 [19], C9 [20], I5, I6, I7, I8 [21]) or transfer learning with synthetic data (C7, C8 [22], I2, I3 [23]), while only two classifiers perform adversarial training using solely the original set (C1, I1 [24]). Some work propose more sophisticated strategies such as: (i) weight averaging (C9 [20], I5, I6, I7, I8 [21]); (ii) ODE networks (C10 [25]); (iii) robust residual blocks (C11 [26]); (iv) modifications on the optimization process (C12 [27], C13 [28], C14 [29], C15 [30]).
@@
Adversarial setup – We implement the under-confidence and over-confidence attacks by employing the cross-entropy loss between the perturbed input and the targets specified in  Sect. 3. We then leverage the Projected Gradient Descent (PGD) method [11] for crafting the adversarial perturbation. When running the attack, we constrain the $l_{\infty}$ norm of the adversarial perturbation with a budget $\epsilon$, that we set as $8/255$ for CIFAR-10 and $4/255$ for ImageNet, accordingly to their original RobustBench evaluation. Lastly, we set the number of attack iterations to 150, observing that this number is suitable as all example losses reached convergence.
@@
Evaluation metrics – In our empirical evaluation, we use the MUS and MSUS measures discussed in  Sect. 3. For the calibration case study, however, a standard calibration measure such as the ECE would capture the miscalibration error but not tell if the model is under or over-confident. Since our study deals with increasing and decreasing the confidence scores of the adversarial samples, it is crucial to use a measure that captures such a phenomenon. Accordingly, we propose a signed version of the Expected Calibration Error, which we refer to as s-ECE (Eq. 19), by taking the signed subtraction between the expected and the observed probabilities (thus, without the absolute value used for the standard ECE).
@@
5.2 Experimental Results
@@
In C1, C3, C4, and C8 (which we denote with group $\diamondsuit$), the mass of the entropy distribution before the attack is primarily concentrated in low-entropy values, with a mode near $\sim 0.1$ and a mean of $\sim 0.6$. The over-confidence attack primarily influences the samples having higher entropies, moving the mass toward the low-entropy mode. The distribution after the under-confidence attack, instead, reveals a bimodal trend: apparently, some of the samples keep lower entropy values, while some others drastically move toward higher entropy regions; such phenomenon generates two modes on the new distribution, one around $\sim 0.1$ and one near $\sim 1.6$. C5, C6, C11, and C13 (group $\clubsuit$) are characterized by a near-uniform distribution before the attack confined between $0.6$ and $1.8$. After the attacks, the distributions gain a soft positive (for the over-confidence attack) and negative (for the under-confidence attack) skewness, shifting the mean of about $0.4$ from its clean value. Even though we cannot group the remaining methods, some present behaviors reported above, such as the bimodal distribution for the under-confidence distribution (C7, C9, C14). C2 and C10 are the best-performing models: in C2, the distribution softly shifts when under attack, while C10 remains almost unchanged. Interestingly, while C10 presents very low clean entropy, C2 has one of the higher clean entropy among the models: such behavior reflects respectively their low and high clean accuracies reported on the RobustBench leaderboard [12].
@@
The clean entropy distributions of I1, I3, and I4 (group $\heartsuit$) present a mode of around $\sim 0.2$; however, a significant part of their mass is also distributed into a considerable interval of medium entropy levels, mainly between $1.0$ and $5.0$. The over-confidence attacks successfully shift a significant part of the mass onto the mode around $\sim 0.2$. After the under-confidence attack, similarly, a substantial mass shift, generating a mode around $\sim 5.2$. The transformer architectures (I5, I6, I7, and I8) (group $\spadesuit$) seem less sensible to both attacks, particularly to over-confidence attacks. Indeed, only I5 presents a mode around higher entropy values, so we expect only I6, I7, and I8 to be generally more robust than the remaining classifiers.
Impact on the model’s calibration – We now show the effect of the over- and under-confidence attacks on the calibration of adversarial trained models. Fig. 4 and Fig. 5 show the reliability diagrams of the 23 robust models for both the analysed data sets, while Table 1 and Table 2 report their s-ECE scores.
@@
Table 1: s-ECE before and after the over- and under-confidence attacks for each CIFAR-10 model.
@@
Table 2: s-ECE before and after the over- and under-confidence attacks for each ImageNet model.
@@
The majority of the clean reliability curves are positioned above the identity line [^2] Note that the spikes in models C1, C3, C10, and C14 do not significantly affect the s-ECE, due to the negligible impact of their small mass on the weighted sum., which empirically supports the findings of Theorem 4.1, i.e., adversarial trained models tend to be under-confident. The “under-confident nature” of adversarial trained models makes them naturally robust against over-confidence attacks (in line with our conjecture in  Sect. 4): indeed, as we can see from Table 1 and Table 2, even after the attack only a few CIFAR-10 models and half of the ImageNet models reach negative s-ECE values (i.e., become over-confident).
Discussion – In LABEL:tab:leaderboard we present a leaderboard ranking models based on their robustness against uncertainty attacks, sorted by MUS.
Table 3: Mean ($\mu$) and standard deviation ($\sigma$) of the MUS and MSUS of each CIFAR-10 and ImageNet model (sorted by $\textbf{MUS}^{\dagger}$). Denoted with * we report potentially unreliable values (see Sect. 5.2).
@@
Engstrom2019Robustness $\heartsuit$ I1 $00.00\%$ $29.22\%$ $2.272$ $1.477$ $7.346$ $7.700$ $6$

### v2

@@ IV Theoretical Analysis on Adversarial Uncertainty Robustness
A recent work [8] likewise raised this observation, but it is still limited to empirically analyzing a small set of models and only against specific attack instances. To this aim, instead of considering specific attack instances, we answer our research question by analyzing the Uncertainty Span of adversarial trained models, i.e., consider both the under- and over-confidence cases, in accordance with  Sect. III, drawing general conclusions on these models’ robustness against any attack instance. Since the predictive entropy of a sample increases when approaching the decision boundary, it is reasonable to think that adversarial training may also protect against under-confidence attacks: in fact, adversarial training forces the model to be robust to label-flips (i.e., to samples crossing the boundary), hindering any potential attacker interested in increasing uncertainty. For over-confidence attacks instead, we cannot draw similar conclusions: indeed, being robust to label flips does not guarantee an entropy reduction since the regions where the entropy is minimized do not lie on the decision boundary.
@@
IV-A On the Convergence of Adversarial Trained Models
@@
Lemma IV.1 represents a well-known result of probabilistic machine learning [15], stating that the optimal classifier should be perfectly calibrated. This phenomenon should be considered when dealing with aleatoric uncertainty because it asserts that lower entropy does not always go along with more accurate classifiers. From this starting point, we will now show that such conditions slightly change in the presence of an attacker.
@@
$\mathcal{L}^{\boldsymbol{\theta}_{R}}(\mathcal{D}_{\textrm{adv}}^{\boldsymbol{x}})$ $=\underbrace{p(\textrm{A}|\boldsymbol{x}+\delta_{\textrm{A}})}_{z}\cdot\underbrace{H(\boldsymbol{t_{A}},f^{\boldsymbol{\theta}}(\boldsymbol{x}+\delta_{\textrm{A}}))}_{-\log(\alpha-\beta_{A})}+$
$+\underbrace{p(\textrm{B}|\boldsymbol{x}+\delta_{\textrm{B}})}_{1-z}\cdot\underbrace{H(\boldsymbol{t_{B}},f^{\boldsymbol{\theta}}(\boldsymbol{x}+\delta_{\textrm{B}}))}_{-\log(1-\alpha-\beta_{B})}$
@@
Contrary to Lemma IV.1, when using adversarial training, the loss does not equal a cross-entropy between two distinct components. Therefore, we need to compute the values for which the partial derivative with respect to $\alpha$ equals zero:
$\implies\frac{\partial\mathcal{L}}{\partial\alpha}=0\implies\frac{1-z}{1-\alpha-\beta}-\frac{z}{\alpha-\beta}=0$
$\implies\alpha=z-\beta\cdot(2z-1)$
IV-B Demystifying the Underconfidence of Robust Models
From Lemma IV.1 and Lemma IV.2, it is possible to derive that, in the presence of an attacker, a classifier inevitably deviates from the trajectory of the optimization process that it would have followed in a standard, clean optimization process. Interestingly, the entropy of an adversarial-trained classifier is higher compared to standard training. We can easily prove this inequality starting from the two lemmas:
Theorem IV.1.
@@ IV-B Demystifying the Underconfidence of Robust Models
From Lemma IV.1, we know $f^{\boldsymbol{\theta}_{S}}(x)=[z,1-z]^{\intercal}$, while from Lemma IV.2 we know that in the presence of an attacker, the optimum deviates from an offset amount of $o=\beta(2z-1)$: $f^{\boldsymbol{\theta}_{R}}(x)=[z-o,1-z+o]^{\intercal}$. Within this interval, the entropy $H(f^{\boldsymbol{\theta}_{R}}(x))$ reaches its minimum when the offset equals zero and, thus, when $f^{\boldsymbol{\theta}_{S}}(x)=f^{\boldsymbol{\theta}_{R}}(x)$; therefore, $H(f^{\boldsymbol{\theta}_{S}}(x))\leq H(f^{\boldsymbol{\theta}_{R}}(x))$. ∎
@@
From Theorem IV.1, we can draw some conclusions on the robustness against uncertainty attacks by studying the effect of adversarial training on the Uncertainty Span. Albeit  Theorem IV.1 proves the tendency of adversarial training to produce underconfident models, it remains factual the robustness of such models against traditional adversarial examples, i.e., crafted to induce a misclassification on a sample [12]. In order to succeed, the adversarial example requires crossing the decision boundary; in a binary classification problem, such conditions correspond precisely to reaching the maximum possible entropy. Consequently, reducing the number of samples for which a traditional adversarial example exists (i.e., increasing the robust accuracy) directly implies reducing the upper bound of the uncertainty span. Concerning the robustness against over-confidence attacks, findings from  Theorem IV.1 suggest greater robustness against such attacks as well. This theorem shows that adversarial-trained models output predictions with higher entropies than undefended models. Suppose we aim to decrease it until reaching a desired value of $v$: in that case, the initially higher entropy of the adversarial trained model makes it more challenging to lower it to $v$, as the starting values are already higher and, therefore, need a significant reduction.
@@
Discussion – In safe environments (without any attacker), a model whose purpose is to solve a classification problem takes advantage of correctly modeling the true posterior $p(y|\boldsymbol{x})$ for any given input sample. Nevertheless, in the presence of an attacker, the classifier tends to balance good generalization capabilities and security guarantees. Such a balance compromises not only the model’s accuracy (as we already know from previous literature) but also comes at the price of sacrificing a certain amount of confidence in the predictions. As a result, adversarial-trained models disclose robust characteristics against uncertainty attacks, providing theoretical support for the claim that adversarial training makes models more resilient to over- and under-confidence attacks.
V Experimental Analysis
@@
We now present the experimental setup employed to empirically validate our theoretical findings by analysing the Uncertainty Span on a selection of 23 state-of-the-art models from the well-known Robustbench repository [13]. In addition, we also consider a concrete case study where an attacker is interested in tampering with the model’s calibration, proving the robustness of adversarial trained models for a real scenario. The code for the reproducibility of the experiments is available at the following link: https://github.com/pralab/UncertaintyAdversarialRobustness
V-A Experimental Setup
@@
Models – For a comprehensive investigation, we consider 15 CIFAR-10 (denoted as C1-C15) and 8 ImageNet (denoted as I1-I8) standard robust classifiers from the Robustbench repository [13], encompassing many different model architectures and robust accuracies, which are summarized in Table I. We consider a wide range of different network architectures, including Convolutional Neural Networks (CNN) and Visual Tranformers (ViT); specifically, for analysing the uncertainty robustness when the model’s size changes, we selected a wide range of model capacity, including ResNet-18, ResNet-50, ResNet-152, WideResNet-28-10, WideResNet-70-10, RaWideResNet-28-10, Swin-B, Swin-L, ConvNeXt-B and ConvNeXt-L. Many of them rely on data augmentation (C2 [18], C3, C4 [19], C5,C6 [19], C9 [20], I5, I6, I7, I8 [21]) or transfer learning with synthetic data (C7, C8 [22], I2, I3 [23]), while only two classifiers perform adversarial training using solely the original set (C1, I1 [24]). Some work proposes more sophisticated strategies such as: (i) weight averaging (C9 [20], I5, I6, I7, I8 [21]); (ii) ODE networks (C10 [25]); (iii) robust residual blocks (C11 [26]); (iv) modifications on the optimization process (C12 [27], C13 [28], C14 [29], C15 [30]).
ID RobustBench Name Architecture
CIFAR-10
C1 [24] Engstrom2019Robustness ResNet-50
C2 [18] Addepalli2022Efficient_RN18 ResNet-18
C3 [19] Gowal2021Improving_28_10_ddpm_100m WideResNet-28-10
C4 [19] Gowal2021Improving_70_16_ddpm_100m WideResNet-70-16
C5 [31] Wang2023Better_WRN-28-10 WideResNet-28-10
C6 [31] Wang2023Better_WRN-70-16 WideResNet-70-16
C7 [22] Sehwag2021Proxy_R18 ResNet-18
C8 [22] Sehwag2021Proxy_ResNest152 ResNet-152
C9 [20] Rebuffi2021Fixing_70_16_cutmix_extra WideResNet-70-16
C10 [25] Kang2021Stable WideResNet-70-16
C11 [26] Peng2023Robust RaWideResNet-70-16
C12 [27] Addepalli2021Towards_RN18 ResNet-18
C13 [28] Cui2023Decoupled_WRN-28-10 WideResNet-28-10
C14 [29] Xu2023Exploring_WRN-28-10 WideResNet-28-10
C15 [30] Pang2022Robustness_WRN70_16 WideResNet-70-16
ImageNet
I1 [24] Engstrom2019Robustness ResNet-50
I2 [23] Salman2020Do_R18 ResNet-18
I3 [23] Salman2020Do_R50 ResNet-50
I4 [32] Wong2020Fast ResNet-50
I5 [21] Liu2023Comprehensive_Swin-B Swin-B
I6 [21] Liu2023Comprehensive_Swin-L Swin-L
I7 [21] Liu2023Comprehensive_ConvNeXt-B ConvNeXt-B
I8 [21] Liu2023Comprehensive_ConvNeXt-L ConvNeXt-L
TABLE I: The RobustBench models used for our analysis. For more details on the specific hyperparameters used during the adversarial training phase, we refer the reader to the original papers.
@@
Adversarial setup – We implement the under-confidence and over-confidence attacks by employing the cross-entropy loss between the perturbed input and the targets specified in  Sect. III. We then leverage the Projected Gradient Descent (PGD) method [12] for crafting the adversarial perturbation. When running the attack, we constrain the $l_{\infty}$ norm of the adversarial perturbation with a budget $\epsilon$, which we set as $8/255$ for CIFAR-10 and $4/255$ for ImageNet, accordingly to their original RobustBench evaluation. Lastly, we set the number of attack iterations to 150, which is slightly higher compared to standard approaches used in literature for traditional evasion attacks targeting the predictions [13, 33]. Moreover, we empirically observed that this is suitable as all example losses reached convergence.
OOD and Open-Set setup – To demonstrate how the analysis of the uncertainty span can be applied to concrete uncertainty attack scenarios, we instantiate our framework in both out-of-distribution (OOD) and Open-Set recognition (OSR) settings. As discussed in Sect. III-A and first proposed in [6], these attacks are designed to disguise samples that are semantically outside the training distribution, making them appear indistinguishable from in-distribution data. To this end, we conduct two types of experiments:
  • OOD scenario: we use 800 samples from the MNIST [34] dataset, which are clearly outside the training distribution;
  • OSR scenario: we select 800 samples from 10 novel classes in the CIFAR-100 dataset. These classes are visually similar to the original 10 CIFAR-10 classes but are not part of the training set, thus simulating an open-set scenario [35].
For each scenario, we construct two datasets composed of 8800 samples (8000 in-distribution and 800 outliers): (i) the unperturbed dataset, and (ii) the dataset in which the outliers are perturbed with an over-confidence attack. Finally, we compute the entropy of all the samples in the dataset, which is used to discern in-distribution samples from the outliers.
@@
Evaluation metrics – In our empirical evaluation, we use the MUS and MSUS measures discussed in  Sect. III. For the calibration case study, however, a standard calibration measure such as the ECE would capture the miscalibration error but not tell if the model is under- or over-confident. Since our study deals with increasing and decreasing the confidence scores of the adversarial samples, it is crucial to use a measure that captures such a phenomenon. Accordingly, we propose a signed version of the Expected Calibration Error, which we refer to as s-ECE (Eq. 19), by taking the signed subtraction between the expected and the observed probabilities (thus, without the absolute value used for the standard ECE).
@@
For the OOD and OSR experiments, we follow common evaluation practices [36, 37] by measuring the AUROC, AUPR (IN/OUT), and FPR95TPR.
V-B Experimental Results
@@
In C1, C3, C4, and C8 (which we denote with group $\diamondsuit$), the mass of the entropy distribution before the attack is primarily concentrated in low-entropy values, with a mode near $\sim 0.1$ and a mean of $\sim 0.6$. The over-confidence attack primarily influences the samples having higher entropies, moving the mass toward the low-entropy mode. The distribution after the under-confidence attack, instead, reveals a bimodal trend: apparently, some of the samples keep lower entropy values, while some others drastically move toward higher entropy regions; such a phenomenon generates two modes on the new distribution, one around $\sim 0.1$ and one near $\sim 1.6$. C5, C6, C11, and C13 (group $\clubsuit$) are characterized by a near-uniform distribution before the attack confined between $0.6$ and $1.8$. After the attacks, the distributions gain a soft positive (for the over-confidence attack) and negative (for the under-confidence attack) skewness, shifting the mean of about $0.4$ from its clean value. Even though we cannot group the remaining methods, some present the behaviors reported above, such as the bimodal distribution for the under-confidence distribution (C7, C9, C14). C2 and C10 are the best-performing models: in C2, the distribution softly shifts when under attack, while C10 remains almost unchanged. Interestingly, while C10 presents very low clean entropy, C2 has one of the highest clean entropies among the models: such behavior reflects their low and high clean accuracies, respectively, reported on the RobustBench leaderboard [13].
@@
The clean entropy distributions of I1, I3, and I4 (group $\heartsuit$) present a mode of around $\sim 0.2$; however, a significant part of their mass is also distributed into a considerable interval of medium entropy levels, mainly between $1.0$ and $5.0$. The over-confidence attacks successfully shift a significant part of the mass onto the mode around $\sim 0.2$. After the under-confidence attack, similarly, a substantial mass shift, generating a mode around $\sim 5.2$. The transformer architectures (I5, I6, I7, and I8) (group $\spadesuit$) seem less sensitive to both attacks, particularly to over-confidence attacks. Indeed, only I5 presents a mode around higher entropy values, so we expect only I6, I7, and I8 to be generally more robust than the remaining classifiers.
Impact on the model’s calibration – We now show the effect of the over- and under-confidence attacks on the calibration of adversarial trained models. Fig. 4 and Fig. 5 show the reliability diagrams of the 23 robust models for both the analysed data sets, while Table II and Table III report their s-ECE scores.
@@
TABLE II: s-ECE before and after the over- and under-confidence attacks for each CIFAR-10 model.
@@
TABLE III: s-ECE before and after the over- and under-confidence attacks for each ImageNet model.
@@
The majority of the clean reliability curves are positioned above the identity line [^2] Note that the spikes in models C1, C3, C10, and C14 do not significantly affect the s-ECE, due to the negligible impact of their small mass on the weighted sum., which empirically supports the findings of Theorem IV.1, i.e., adversarial trained models tend to be under-confident. The “under-confident nature” of adversarial trained models makes them naturally robust against over-confidence attacks (in line with our conjecture in  Sect. IV): indeed, as we can see from Table II and Table III, even after the attack only a few CIFAR-10 models and half of the ImageNet models reach negative s-ECE values (i.e., become over-confident).
Discussion – In LABEL:tab:leaderboard we present a leaderboard ranking models based on their robustness against uncertainty attacks, sorted by MUS. For each of the reported values we also show the standard deviation computed on the MUS and MSUS distributions, for strengthen the statistical significance of the results.
TABLE IV: Mean ($\mu$) and standard deviation ($\sigma$) of the MUS and MSUS of each CIFAR-10 and ImageNet model (sorted by $\textbf{MUS}^{\dagger}$), along with their accuracies and RobustBench-based rank. Denoted with * we report potentially unreliable values, as discussed in Sect. V.
@@
Engstrom2019Robustness $\heartsuit$ I1 [24] $62.56\%$ $29.22\%$ $2.272$ $1.477$ $7.346$ $7.700$ $6$